OpenStudy (pythagoras123):

There are 72 students in Grade 6. Each of them paid an equal amount of money for their textbooks. The total amount of money collected is $█35.0█, where two of the digits █ cannot be recognized. Nevertheless, find the amount of money each student paid for the books.

OpenStudy (anonymous):

are they supposed to be the same digit?

OpenStudy (pythagoras123):

No. they represent different digits.

OpenStudy (anonymous):

stumped. are you supposed to solve congruences?

OpenStudy (pythagoras123):

It has something to do with divisibility rules, I think...

OpenStudy (anonymous):

135.00/72=1.875 235.00/72=3.2638..... 335/72=4.65....

OpenStudy (anonymous):

only the first one is excat value, but ofcourse I assumed .00

OpenStudy (anonymous):

you sure they are not the same number..

OpenStudy (pythagoras123):

Yes. As a facilitation, the total amount paid is $635.04, and each student paid $8.82. But I don't know how to arrive at the answer

OpenStudy (anonymous):

how did you find total amount which is 635.04

OpenStudy (anonymous):

635.04/72=8.82